OpenStudy (anonymous):

using distance formula (1, 1) and (-4, 1) A. 4 B. SQRT(61) C.SQRT(290) D. 5 E. SQRT(173) F. SQRT(180)

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) what you got ?

hartnn (hartnn):

as x1-x2= ... ? and y1-y2=... ?

OpenStudy (anonymous):

(-4 - (1)) sqrt 2+(1 - 1)) sqrt 2

hartnn (hartnn):

and 1-1 =0 what is -4-1 =... ?

OpenStudy (anonymous):

(-4 + 1) sqrt 2 ?

hartnn (hartnn):

no, \(\sqrt{(-4-1})^2=\sqrt{5^2}=.. ? \)

hartnn (hartnn):

not getting ?

OpenStudy (anonymous):

not really, Ive been doing math all day so my brains hurting lol

hartnn (hartnn):

-4-1 =-5 so, \(\sqrt{(-5)^2}=\sqrt{(5)^2}=\sqrt{25}=5\) got this ?

OpenStudy (anonymous):

ohh so i was thinking right

hartnn (hartnn):

yes :)

OpenStudy (anonymous):

yay :D